Find an equation of the the line satisfying the given conditions.
Through (3, 15); parallel to 2.5x - 0.9y = -0.6

Answer by checkley77(12844)

Line equation is Y=mX+b
Through (3, 15); parallel to 2.5x - 0.9y = -0.6
-.9y=-2.5x-.6
.9y=2.5x+.6
A parallel line has the same slope(m).
15=2.5*3+b
15=7.5+b
b=12-7.5
b=4.5 ans. for the y intercept.
y=-2.5x+4.5 ans. for this line equation through (3,15)